# Spectral properties from an efficient analytical representation of the $GW$ self-energy within a multipole approximation

### v1

@@ III.3 The quasiparticle picture in MPA models
In both the cases of $\mathrm{Im}~{}Z=0$ and $\mathrm{Im}~{}Z=0.3$, as $\mathrm{Re}~{}Z$ decreases, $\mathrm{Re}~{}[\epsilon^{\text{s1}}_{1}]$ goes to negative values with decreasing spectral weight $\mathrm{Re}~{}[P^{\text{s1}}_{1}]$, while $\mathrm{Re}~{}[\epsilon^{\text{s1}}_{2}]$ increases from $\xi$ (solid gray line in panel (a)), with increasing $\mathrm{Re}~{}[P^{\text{s1}}_{2}]$. The finite $\mathrm{Im}~{}Z$ induces a finite imaginary part in the poles and residues. As $\mathrm{Re}~{}Z$ decreases, both $\mathrm{Im}~{}[\epsilon^{\text{s1}}_{p}]$ increase in modulus, while $\mathrm{Im}~{}[P^{\text{s1}}_{p}]$ decrease despite $\mathrm{Im}~{}Z$ being constant. The broadening of the poles avoids the infinite correlation limit affecting the curvature of $\mathrm{Re}~{}[\epsilon^{\text{s1}}_{p}]$ and $\mathrm{Re}~{}[P^{\text{s1}}_{p}]$ (red vs. blue curves in panels (a, b)).
@@
For all values of $Z$, the first pole has the largest spectral weight and thus $Z^{\text{MPA-}G}_{\text{s1}}=P^{\text{s1}}_{1}$. As illustrated by the dotted black line in panel (b), when we move from the independent-particle limit, the renormalization factor $Z_{\text{s1}}^{\text{lin}}$ starts deviating from $Z_{\text{s1}}^{\text{MPA-}G}$ (dark blue curve), while for $\mathrm{Im}~{}Z=0.3$ (dark red curve), both definitions already differ at $\mathrm{Re}~{}Z=1$ and only their real parts coincide around $\mathrm{Re}~{}[Z]=0.8$. The plot shows that, if $Z$ is taken as an indicator of the validity of the QP picture, the two definitions give inconsistent results, since $\mathrm{Re}~{}[Z_{\text{s1}}^{\text{MPA-}G}]>0.5$ in the whole interval, even when $\mathrm{Re}~{}[Z_{\text{s1}}^{\text{lin}}]<0.5$. Therefore, the local definition of the renormalization factor can only be used in the regime of weak correlation ($Z\approx 1$), while in the large correlation limit ($Z\to 0$), the energy position of the poles diverges ($\mathrm{Re}~{}[\epsilon^{\text{s1}}_{p}]\to\mp\infty$) with similar spectral weight ($\mathrm{Re}~{}[P^{\text{s1}}_{p}]\to 0.5$).
@@
In Fig. 9 we analyze the effects of the static term. The plots are analogous to Fig. 8. Notice that we have extended the range of the plots to $\mathrm{Re}~{}Z=-0.1$, to test the QP picture even for extreme values. In this case, we have fixed $\mathrm{Im}~{}Z=0.01$ and considered three values of $x$, resulting in a picture similar to the one described in Fig. 8, where $\mathrm{Re}~{}[P^{\text{s1}}_{1}]$ is always larger than $0.5$. The value $x=0$ corresponds to the previously discussed case of a self-energy with only the correlation term. The main effect of a finite value, illustrated with $x=\pm 0.3$, is to change the concavity of the residues according to its sign (panels (b-c)). At variance with $x\leq 0$, for positive values $\mathrm{Re}~{}[P^{\text{s1}}_{1}]<\mathrm{Re}~{}Z$ close to $\mathrm{Re}~{}Z\to 1$, therefore, in this region Eq. (16) produces lower values than the renormalization factor of Eq. (7).
The toy model $\Sigma^{\text{s1}}$ effectively explains the possible causes of the differences between the renormalization factors computed in Fig. 4 with Eqs. (7) and (16). Additional variations in these parameters due to a richer structure of the self-energy are explained by the two-pole self-energy model $\Sigma^{\text{s2}}$ presented in Appendix D. However, if $\Sigma$ is dominated by a single pole the overall picture is the same. Therefore, $\Sigma^{\text{s1}}$ captures the most typical situations when solving the QP equation in prototypical materials. Similar toy MPA models could help clarifying situations in which the QP picture does not hold [4, 64].
IV Conclusions
We have presented MPA-$\Sigma$, a robust method to efficiently approximate the frequency dependence of the $GW$ self-energy as a multipole-Padé representation, typically with around 10 poles. The method is similar to the multipole approximation for the screening interaction, MPA-$W$. Analogous to MPA-$W$, MPA-$\Sigma$ is built from numerical data evaluated on around 20 frequency points in the complex plane, thus avoiding explicit evaluations of the self-energy on dense frequency grids (of the order of 1000 frequencies in our cases). Using MPA-$\Sigma$, it is also possible to solve the QP equation analytically and obtain an MPA-$G$ representation of the interacting Green’s function, from which all the spectral properties can be easily extracted, including the position, broadening, and residue of the QP pole and its satellites.
Combining MPA-$W$ and MPA-$\Sigma$ is a computationally powerful approach, reducing the number of frequency evaluations by a factor $\sim 50^{2}$, compared to full-frequency approaches, while providing spectra with comparable numerical accuracy. The excellent accuracy of the novel method has been verified for several materials: bulk Si, Na and Cu, monolayer MoS2, the NaCl ion-pair and the F2 molecule. The efficiency of the MPA method allows us to compute $\Sigma$ and $G$ spectra in a wide energy range. In particular, our results for NaCl and F2 exhibit features beyond the typical energy range of the $GW$ spectra found in the literature for such molecular species. We have also presented a method for interpolating the spectra of higher-dimensional systems in momentum space, and construct $\Sigma$ and $G$ spectral band structures. For Na, Si, monolayer MoS2, and Cu, we report full $\Sigma$ and $G$ spectral band structures, while isolating the contributions from the QP pole and the satellites.
The spectral weights of both the QP pole and the satellites computed with MPA-$G$ comply with the sum rule for the number of particles/holes, providing a more general way to define the renormalization factor. Therefore, such analytical representations can be useful in understanding the physical nature of the QP picture and the renormalization factor. We have presented a toy model that captures the most typical situations when solving the QP equation in prototypical materials, exposing the limitations of the renormalization factor as defined from the local derivative of the self-energy, to describe the spectral weight of the most intense peak of the Green’s function beyond the weak correlation regime. Our findings suggest revisiting some literature results on the renormalization factor. They also open the possibility to expand this kind of analysis to the search of materials with exotic spectral properties.
Acknowledgments
This work was funded by the Research Council of Norway through the MORTY project (315330). Access to high performance computing resources was provided by UNINETT Sigma2 (NN9711K) in Norway, and by EuroHPC Joint Undertaking through the project EHPC-EXT-2022E01-022 that grants access to Leonardo-Booster@Cineca, Italy. We acknowledge stimulating discussions with Andrea Ferretti, Kristian S. Thygesen and Mikael Kuisma.
Appendix A MPA-$\Sigma$ interpolation
Consider the exchange and correlation terms of the self-energy, $\Sigma=\Sigma_{\text{x}}+\Sigma_{\text{c}}$; for any state $n{\mathbf{k}}$, the multipole-Padé representation of $\Sigma_{\text{c}}$ can be obtained by solving the following system of $2n_{\Sigma}$ equations and variables:
$\Sigma_{\text{c}}^{\text{MPA}}(z_{i})\equiv\sum_{p}^{n_{\Sigma}}\frac{S_{p}}{z_{i}-\xi_{p}}=\Sigma_{c}(z_{i})\text{, }i=1,\ldots,2n_{\Sigma}.$ (25)
Since this model uses an interpolation to reconstruct $\Sigma_{\text{c}}(z)$, the solution will depend on the selected sampling points, $\{z_{i}\}$.
To solve Eq. (25), we follow a similar procedure to the one used for $W_{c}$, by rewriting $\Sigma_{\text{c}}^{\text{MPA}}(z)$ in its Padé form, i.e. as a fraction of two polynomials:
$\sum_{p}^{n_{\Sigma}}\frac{S_{p}}{z-\xi_{p}}=\frac{A_{n_{\Sigma}-1}(z)}{B_{n_{\Sigma}}(z)},$ (26)
where the coefficients of the $B_{n_{\Sigma}}(z)$ polynomial can be evaluated from the numerical reference data, $\{z_{i},\Sigma_{c}(z_{i})\}$, using one of the two methods developed in Ref. [39], based on linear algebra and Thiele’s Padé interpolation (see details in section I of Ref. [79]). Moreover, its factorization can be performed using the Companion matrix method [39] and is given by
$B_{n_{\Sigma}}(z)=\prod_{p}^{n_{\Sigma}}(z-\xi_{p}).$ (27)
In both methods the sampling points are divided in two sets, which helps to separate the problem of finding the poles, $\xi_{p}$, from the much simpler problem of finding the residues, $S_{p}$, once the poles are known. Such separation is computational advantageous since the nonlinear problem of $2n_{\Sigma}$ variables in Eq. (25), is reduced to two problems of size $n_{\Sigma}$, one nonlinear for the poles and the other linear for the residues. Moreover, by first obtaining the poles, it is then possible to apply physical constrains. We impose to $\xi_{p}=\zeta_{p}+i\varsigma_{p}$ that $\zeta_{p}/\varsigma_{p}<-1$, which guarantees the time ordering and that the poles lay closer to the real frequency axis, as was done for MPA-$W$ [39]. The residues can then be found by solving a simple linear least squares problem (see details in section I of Ref. [79]).
@@ Appendix A MPA-$\Sigma$ interpolation
Figure 10: Real part of the poles and residues of the Green’s function model of Eq. (38) as a function of $\mathrm{Re}~{}Z$, where we have fixed $\mathrm{Im}~{}Z=0$.
Appendix B MPA-$G$ from MPA-$\Sigma$
Given the MPA-$\Sigma$ representation of the correlation self-energy with $n_{\Sigma}$ poles, the total self-energy can be written in its Padé representation as
@@ Appendix B MPA-$G$ from MPA-$\Sigma$
$\Sigma^{\text{MPA}}(z)\equiv\Sigma_{x}+\Sigma_{c}^{\text{MPA}}(z)=\frac{\Sigma_{x}B_{n_{\Sigma}}(z)+A_{n_{\Sigma}-1}(z)}{B_{n_{\Sigma}}(z)}.$ (28)
Notice that if we apply physical constrains after the factorization in Eq. (27), and fit the residues thereafter, we will need to reconstruct both the $A_{n_{\Sigma}-1}(z)$ and $B_{n_{\Sigma}}(z)$ polynomials from the new poles and residues using Eq. (28), which is straightforward.
We can then compute the interacting Green’s function of Eq. (3) in the MPA approach as
$G^{\text{MPA}}(z)$ $\equiv\frac{B_{n_{\Sigma}}(z)}{C_{n_{\Sigma}+1}(z)}$ (29)
$C_{n_{\Sigma}+1}(z)$ $=zB_{n_{\Sigma}}(z)-\Sigma_{x}B_{n_{\Sigma}}(z)-A_{n_{\Sigma}-1}(z),$
The poles of $G$ will then correspond to the zeros of $C_{n_{\Sigma}+1}(z)$. Analogous to $B_{n_{\Sigma}}(z)$, the $C_{n_{\Sigma}+1}(z)$ polynomial can be factorized, e.g. using the Companion matrix method [39]:
$C_{n_{\Sigma}+1}(z)=\prod_{p}^{n_{\Sigma}+1}(z-\epsilon_{p}),$ (30)
while the residues can be computed using the residue theorem (see also Ref. [57]), as
$P_{p}=\frac{\prod_{i}^{n_{\Sigma}}(\epsilon_{p}-\xi_{i})}{\prod_{i\neq p}^{n_{\Sigma}+1}(\epsilon_{p}-\epsilon_{i})},$ (31)
from where it follows that the sum rule of Eq. (15) is satisfied. The resulting multipole-Padé representation of $G$ is then:
$G^{\text{MPA}}(z)=\sum_{p=1}^{n_{\Sigma}+1}\frac{P_{p}}{z-\epsilon_{p}}.$ (32)
Appendix C Spline interpolation in ${\mathbf{k}}$-space
@@
To obtain smooth plots of the $\Sigma$ and $G$ spectral band structures, similar to the band structures of the response function computed in Ref. [80], we need to perform an interpolation on ${\mathbf{k}}$ space for each frequency $z_{i}$. Such an interpolation can be cumbersome due to the dispersion of the different bands. The MPA-$\Sigma$ representation can simplify this interpolation since it is sufficient to interpolate the poles and the residues of the multipole-Padé model. The interpolation of the numerical data can be also simplified by considering an auxiliary set of frequencies centered at the KS energies, $z^{\prime}_{i}\equiv z_{i}-\epsilon^{\text{KS}}_{n{\mathbf{k}}}$. Since $z^{\prime}_{i}$ carries the dispersion of the bands, $\Sigma(z^{\prime}_{i},{\mathbf{k}})$ and $G(z^{\prime}_{i},{\mathbf{k}})$ are much smoother than the original $\Sigma(z_{i},{\mathbf{k}})$ and $G(z_{i},{\mathbf{k}})$.
@@
where $f^{\Sigma}_{n}$ and $f^{G}_{n}$ are the interpolating function of $\Sigma$ and $G$ respectively. The interpolation for the $\{z_{i}\}$ frequencies can then be obtained as
@@
which can be used to evaluate the final spectra on a much denser ${\mathbf{k}}$-grid. We have used first-order splines as the interpolating functions, which is sufficient to obtain smooth spectra for our calculations, although this approach can be applied with other interpolating functions as well.
Appendix D Toy MPA-$\Sigma$ model with two poles
Similar to the $\Sigma^{\text{s1}}$ model discussed in Sec. III.3, here we introduce a toy MPA-$\Sigma$ model with two poles, one at $\epsilon_{1}=\xi$ and the other at a larger energy $\epsilon_{2}=(1+a)\xi$:
$\Sigma^{\text{s2}}(\omega)=x\xi+\\ \frac{1-2Z}{2Z}\frac{\xi^{2}}{\omega-\xi}+\frac{(1+a)^{2}}{2Z}\frac{\xi^{2}}{\omega+(1+a)\xi}.$ (36)
where the residues, $S_{1}=\xi^{2}(1-2Z)/(2Z)$ and $S_{2}=\xi^{2}(1+a)^{2}/(2Z)$ are constrained so that the following expression remains invariant, as for $\Sigma^{\text{s1}}$:
$Z^{\text{lin}}_{\text{s2}}\equiv\left[1-\frac{\partial\Sigma^{\text{s2}}(\omega)}{\partial\omega}\bigg{|}_{\omega=0}\right]^{-1}=Z.$ (37)
$G^{\text{s2}}(\omega)=\sum_{p=1}^{3}\frac{P^{\text{s2}}_{p}(x,Z,a)}{\omega-\epsilon^{\text{s2}}_{p}(x,Z,a,\xi)}.$ (38)
Notice that for $a=0$, $\Sigma^{\text{s2}}$ simplifies to $\Sigma^{\text{s1}}$, and therefore this parameter can be used to turn on the second pole. We then fix $x=0$ and $\mathrm{Im}~{}Z=0$, and compare the two models. Fig. 10 shows the real part of the scaled poles, $\epsilon^{\text{s2}}_{p}/\xi$ (panel (a)), and the residues, $P_{p}^{\text{s2}}$ (panel (b)), of $G^{\text{s2}}$ as functions of $\mathrm{Re}~{}Z$, for $a=0$ (blue curves) and $a=1$ (orange curves). The overall picture discussed in Sec. III.3 is similar for the two models, with the additional third pole remaining around $\epsilon^{\text{s2}}_{3}=\xi$ with a vanishing residue $P^{\text{s2}}_{3}$. However, a finite $a$ induces a finite $P^{\text{s2}}_{3}$, with its larger modulus at $\mathrm{Re}~{}Z=1$, which introduces a deviation between $Z^{\text{lin}}_{\text{s2}}$ and $Z^{\text{MPA-}G}_{\text{s2}}=P^{\text{s2}}_{1}$ around $Z=1$.

### v2

@@ A.1 Toy MPA-$\Sigma$ model with a single pole
In both the cases of $\mathrm{Im}~{}\lambda=0$ and $0.3$, as $\mathrm{Re}~{}\lambda$ increases, $\mathrm{Re}[\epsilon^{\text{s1}}_{1}]$ goes to negative values with decreasing spectral weight $\mathrm{Re}[Z^{\text{s1}}_{1}]$, while $\mathrm{Re}[\epsilon^{\text{s1}}_{2}]$ increases from $\xi$ (solid gray line in panel (a)), with increasing $\mathrm{Re}[Z^{\text{s1}}_{2}]$. The finite $\mathrm{Im}~{}\lambda$ induces a finite imaginary part in the poles and residues. As $\mathrm{Re}~{}\lambda$ increases, both $\mathrm{Im}[\epsilon^{\text{s1}}_{p}]$ increase in modulus, while $\mathrm{Im}[Z^{\text{s1}}_{p}]$ decrease despite $\mathrm{Im}~{}\lambda$ being constant. The broadening of the poles avoids the infinite correlation limit, affecting the curvature of $\mathrm{Re}[\epsilon^{\text{s1}}_{p}]$ and $\mathrm{Re}[Z^{\text{s1}}_{p}]$ (orange vs. blue curves in panels (a, b)).
@@
As illustrated by the dotted black line in Fig. 7(b), when we move from the independent-particle limit, the renormalization factor $Z_{\text{s1}}^{\text{lin}}$ starts deviating from $Z_{\text{s1}}^{\text{MPA-}G}$ (dark blue curve), while for $\mathrm{Im}~{}\lambda=0.3$ (dark orange curve), both definitions already differ at $\mathrm{Re}~{}\lambda=0$ and only their real parts coincide around $\mathrm{Re}~{}\lambda=0.2$. The plot shows that, $Z^{\text{lin}}$ cannot be taken as an indicator of the validity of the QP picture beyond the regime of weak correlation ($\lambda\approx 0$), since $\mathrm{Re}[Z_{\text{s1}}^{\text{MPA-}G}]>0.5$ in the whole interval, even when $\mathrm{Re}[Z_{\text{s1}}^{\text{lin}}]<0.5$. In the large correlation limit ($\lambda\to 1$), the energy position of the poles diverges ($\mathrm{Re}[\epsilon^{\text{s1}}_{1,2}]\to\mp\infty$) with similar spectral weight ($\mathrm{Re}[Z^{\text{s1}}_{1,2}]\to 0.5$).
@@
In Fig. 8 we analyze the effects of the static term. The plots are analogous to Fig. 7. Notice that we have extended the range of the plots to $\mathrm{Re}~{}\lambda=1.1$, to test the QP picture even for extreme values. In this case, we have fixed $\mathrm{Im}~{}\lambda=0.01$ and considered three values of $x$, resulting in a picture similar to the one described in Fig. 7, where $\mathrm{Re}[Z^{\text{s1}}_{1}]$ is always larger than $0.5$. The value $x=0$ corresponds to the previously discussed case of a self-energy with only the correlation term. The main effect of a finite value, illustrated with $x=\pm 0.3$, is to change the concavity of the residues according to its sign (orange/green vs. blue curves in panel (b)). At variance with $x\leq 0$, for positive values $\mathrm{Re}[Z_{\text{s1}}^{\text{MPA-}G}]<\mathrm{Re}[Z_{\text{s1}}^{\text{lin}}]$ close to $\mathrm{Re}~{}\lambda\to 0$. In the case of $x>1$, the spectral weight of the satellite is larger than the QP pole, as $\mathrm{Re}[Z^{\text{s1}}_{1}]<\mathrm{Re}[Z^{\text{s1}}_{2}]$.
@@
Figure 9: Real part of the poles (a) and residues (b) of the Green’s function model of Eq. (32) as a function of $\mathrm{Re}~{}\lambda$, where we have fixed $\mathrm{Im}~{}\lambda=0$.
A.2 Toy MPA-$\Sigma$ model with two poles
Similarly to the $\Sigma^{\text{s1}}$ model discussed in Sec. A, here we introduce a toy MPA-$\Sigma$ model with two poles, one at $\epsilon_{1}=\xi$ and the other at a larger energy $\epsilon_{2}=(1+a)\xi$:
$\Sigma^{\text{s2}}(\omega)=x\xi+\\ \frac{2\lambda-1}{2(1-\lambda)}\frac{\xi^{2}}{\omega-\xi}+\frac{(1+a)^{2}}{2(1-\lambda)}\frac{\xi^{2}}{\omega-(1+a)\xi}.$ (30)
where the residues, $S_{1}=\xi^{2}(2\lambda-1)/[2(1-\lambda)]$ and $S_{2}=\xi^{2}(1+a)^{2}/[2(1-\lambda)]$ are constrained so that the following expression remains invariant, as for $\Sigma^{\text{s1}}$:
@@ A.2 Toy MPA-$\Sigma$ model with two poles
$Z^{\text{lin}}_{\text{s2}}\equiv\left[1-\frac{\partial\Sigma^{\text{s2}}(\omega)}{\partial\omega}\bigg{|}_{\omega=0}\right]^{-1}=1-\lambda.$ (31)
$G^{\text{s2}}(\omega)=\sum_{p=1}^{3}\frac{Z^{\text{s2}}_{p}(x,\lambda,a)}{\omega-\epsilon^{\text{s2}}_{p}(x,\lambda,a,\xi)}.$ (32)
Notice that for $a=0$, $\Sigma^{\text{s2}}$ simplifies to $\Sigma^{\text{s1}}$, and therefore this parameter can be used to turn on the second pole. We then fix $x=0$ and $\mathrm{Im}~{}\lambda=0$, and compare the two models. Figure 9 shows the real part of the scaled poles, $\epsilon^{\text{s2}}_{p}/\xi$ [Fig. 9(a)], and the residues, $Z_{p}^{\text{s2}}$ [Fig. 9(b)], of $G^{\text{s2}}$ as functions of $\mathrm{Re}~{}\lambda$, for $a=0$ (blue curves) and $a=1$ (orange curves). The overall picture discussed in Sec. A is similar for the two models, the first one having a satellite increasing from $\epsilon^{\text{s1}}_{2}=\xi$, while the second has two, one increasing from $\epsilon^{\text{s1}}_{2}=(1+a)\xi$ and the second remaining around $\epsilon^{\text{s2}}_{3}=\xi$ with a vanishing residue $Z^{\text{s2}}_{3}$. However, a finite $a$ induces a finite $Z^{\text{s2}}_{3}$, with its larger modulus at $\mathrm{Re}~{}\lambda=0$, which introduces a deviation between $Z^{\text{lin}}_{\text{s2}}$ and $Z^{\text{MPA-}G}_{\text{s2}}$ around $\lambda=0$. Notice that even if $Z^{\text{lin}}_{\text{s2}}=1$ for $\lambda=0$, the linearized QP correction does not vanish for finite $a$, as $\Sigma^{\text{s2}}(\omega=0)/\xi=-(2l+a)/[2(1-l)]$. The weak correlation limit is found for both $\lambda\to 0$ and $a\to 0$ simultaneously.
Appendix B Spline interpolation in ${\mathbf{k}}$-space
@@
To obtain smooth plots of the $\Sigma$ and $G$ spectral band structures, similar to the band structures of the response function computed in Ref. [86], we need to perform an interpolation on the ${\mathbf{k}}$ space for each frequency $z_{i}$. Such an interpolation can be cumbersome due to the dispersion of the different bands. The MPA-$\Sigma$ and MPA-$G$ representations can simplify this interpolation, since it is sufficient to interpolate the poles and the residues of each multipole-Padé model. The interpolation of numerical data can also be simplified by considering an auxiliary set of frequencies centered on the KS energies $z^{\prime}_{i}\equiv z_{i}-\epsilon^{\text{KS}}_{n{\mathbf{k}}}$. Since $z^{\prime}_{i}$ carries the dispersion of the bands, $\Sigma(z^{\prime}_{i},{\mathbf{k}})$ and $G(z^{\prime}_{i},{\mathbf{k}})$ are much smoother than the original $\Sigma(z_{i},{\mathbf{k}})$ and $G(z_{i},{\mathbf{k}})$.
@@
where $f^{\Sigma}_{n}$ and $f^{G}_{n}$ are the interpolating functions of $\Sigma$ and $G$, respectively. The interpolation for the $\{z_{i}\}$ frequencies can then be obtained as
@@
which can be used to evaluate the final spectra on a much denser ${\mathbf{k}}$-grid. We have used first-order splines as interpolating functions, which is sufficient to obtain smooth spectra for our calculations, although this approach can be applied with other interpolating functions as well.